• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Niedziela

    Log entry time 05:20:38 on July28,2003

    Entry number 107614

    keyword=Holding Field rotated to zero

    The target holding field was rotated to zero degrees @4:35, before the NMR/EPR measurements were done.

    4:33 - Laser 1 off
    4:34 - Laser 2 off
    4:35 - Laser 3 off
    4:35 - Begin rotating holding field to 0 from 270
    4:38 - Rotation complete.
    4:38 - Laser 4 on
    4:39 - Laser 6 on
    4:40 - Laser 7 on.

    Waited for temperature to stabilize from 4:40 - 4:50 before conducting polarization tests.